# On the Theory of Risk-Aware Agents: Bridging Actor-Critic and Economics

### v2

@@ C.1 Pessimistic Actor
The pessimistic actor, denoted as $\pi^{p}_{\phi}$, employs a soft policy target for optimization, as outlined in Equation 17. This actor operates entirely off-policy, utilizing data exclusively from the optimistic actor. The soft policy approach not only promotes state-dependent exploration (Haarnoja et al., 2018) but also helps in maintaining the hyperbolic tangent (Tanh) output in a non-saturated state (Wang et al., 2020a). Moreover, the pessimistic actor’s inherent non-zero variance contributes to stabilizing the TD learning for the critic. While separating exploration and exploitation theoretically allows for zero variance in TD target sampling, our empirical evidence suggests that introducing noise regularizes the critic (see Table 1).
[…]
@@ C.2 Optimistic Actor
[…]
@@
    Reducing the degree of off-policy learning - Given that exploration is conducted solely by the optimistic actor, the pessimistic actor updates entirely off-policy, using transitions sampled from the optimistic policy. This can potentially lead to unstable learning, a phenomenon known as the ”deadly triad” (Sutton & Barto, 2018). Integrating KL divergence into the optimistic objective aligns the transitions sampled from the optimistic policy with the expectations under the pessimistic policy, thereby mitigating the extent of off-policy learning.
@@
C.3 Adjustment of $\beta^{o}$ and $\tau$
@@
Since values of $Q^{\mu}_{\theta}$ and $Q^{\sigma}_{\theta}$ depend on reward scales, as well as aleatoric and epistemic uncertainty of the environment, the value of $\beta^{o}$ cannot be easily set. DAC leverages an observation that for $\beta^{o}=-\beta^{p}$ the optimistic actor recovers the objective of the pessimistic actor. Then, $\beta^{o}$ can be defined such that the divergence between the pessimistic baseline policy and the optimistic policy reaches a desired level, as shown in Equation 20. This adaptive approach, as illustrated in Figure 10, accommodates different scales of Q-values and contrasts with setups like OAC, where optimism is predefined by fixing $\beta^{o}$ at a specific value. We find that the optimism and KL weight used by DAC are highly variable between tasks and even phase of the training, showing the advantage of the automatic adjustment mechanism as opposed to using predefined levels. Furthermore, we find that DAC generally uses much higher levels of optimism than values that were found optimal for OAC (Ciosek et al., 2019).
Figure 10: The observed levels of optimism ($\beta^{o}$) and KL weight ($\tau$) during the training in both low and high replay regimes. As follows from the conducted measurements, the levels of optimism used by DAC depend not only upon the tested task, but also upon the phase of the training, with higher levels of optism seen in the latter training phases. $Y$-axis denotes the mean optimism (left axis) and KL weight (right axis) value with 3 standard deviation CI. 10 seeds per task.
C.4 Replay Ratio Scaling
Increasing the replay ratio was the key ingredient in increasing the sample efficiency and final performance of many recent RL agents (Janner et al., 2019; Chen et al., 2020; Hiraoka et al., 2021; Nikishin et al., 2022; Schwarzer et al., 2023). Moreover, many of the novelties brought by those agents were focused on stabilizing the learning given the high-replay regime (Janner et al., 2019; Chen et al., 2020; Hiraoka et al., 2021). A particularly interesting technique is scaling by resetting, as it is applicable to most off-policy RL algorithms. DAC improvements are largely orthogonal to the direction of increasing the replay ratio. Figures 1, 4 and 6 show DAC performance with and without scaling by resetting.
C.5 Non-soft value DAC derivation
In the main body of the paper, we derive and implement DAC as a pure MaxEnt algorithm, such that employs both soft policy and soft values. In this section, we discuss simpler derivations of the policy evaluation and improvement steps in DAC stemming from non-soft value functions. In such setup, value is defined as:
$V^{\pi}(s)=\underset{a\sim\pi}{\mathrm{E}}r_{s,a}+V^{\pi}(s^{\prime})$ (45)
Then, the optimistic actor can be defined as:
$\begin{split}\pi_{\eta}^{o}&=\underset{\pi\in\Pi}{\arg\max}~{}\mathrm{E}~{}e^{~{}2\beta^{o}V^{i}_{\theta}(s)+\tau KL(\pi_{\eta}^{o}(s)|\pi_{\phi}^{p}(s))}\\ &\approx\underset{\pi\in\Pi}{\arg\max}~{}\underset{s\sim\mathcal{D}}{\mathrm{E}}~{}V_{\theta}^{\mu}(s)+\beta^{o}V_{\theta}^{\sigma}(s)-\tau KL(\pi_{\eta}^{o}(s)|\pi_{\phi}^{p}(s))\\ &\approx\underset{\pi\in\Pi}{\arg\max}~{}\underset{s\sim\mathcal{D}}{\mathrm{E}}Q_{\theta}^{\mu}(s,a)+\beta^{o}Q_{\theta}^{\sigma}(s,a)-\tau KL(\pi_{\eta}^{o}(s)|\pi_{\phi}^{p}(s)),~{}~{}~{}~{}a\sim\pi_{\eta}^{o}(s)\end{split}$ (46)
Given a non-soft value definition, the pessimistic actor can be defined as a ”twin” of the optimistic one:
$\begin{split}\pi_{\phi}^{p}&=\underset{\pi\in\Pi}{\arg\max}~{}\underset{s\sim\mathcal{D}}{\mathrm{E}}~{}e^{~{}2\beta^{p}V^{i}_{\theta}(s)+\alpha KL(\pi_{\phi}^{p}|U)}\end{split}$ (47)
Above, $U$ denotes a uniform distribution with its domain over the entire action space. As such, whereas the optimistic actor minimizes the KL divergence to the pessimistic actor, and the pessimistic actor minimizes the KL divergence to a uniform distribution. Such definition retrieves entropy maximizing policy:
$\begin{split}\pi_{\phi}^{p}&\approx\underset{\pi\in\Pi}{\arg\max}~{}\underset{s\sim\mathcal{D}}{\mathrm{E}}~{}V_{\theta}^{\mu}(s)+\beta^{p}V_{\theta}^{\sigma}(s)-KL(\pi_{\phi}^{p}(s)|U(s))\\ &=\underset{\pi\in\Pi}{\arg\max}~{}\underset{s\sim\mathcal{D}}{\mathrm{E}}~{}V_{\theta}^{\mu}(s)+\beta^{p}V_{\theta}^{\sigma}(s)+\alpha\mathcal{H}(s)\\ &\approx\underset{\pi\in\Pi}{\arg\max}~{}\underset{s\sim\mathcal{D}}{\mathrm{E}}~{}Q_{\theta}^{\mu}(s,a)+\beta^{p}Q_{\theta}^{\sigma}(s,a)-\alpha\log\pi_{\phi}^{p}(a|s),~{}~{}~{}~{}a\sim\pi_{\phi}^{p}(s)\end{split}$ (48)
@@ Appendix D Additional Experiments
D.1 Simplified and Regular DAC
In our investigation, we assess the effectiveness of the regular Dual Actor-Critic (DAC) architecture, where the optimistic policy is a perturbation of the pessimistic one, against a simplified architecture that treats the optimistic policy as independent. The primary focus is on performance disparities across all $21$ tasks listed in Table 2, considering both low and high replay ratios. These findings are illustrated in Figure 11.
Figure 11: We compare the performance of the simplified and regular DAC architectures. We find that the simplified architecture yields very similar performance, except for the more complex tasks where the regular architecture dominates. 21 tasks and 10 seeds per task.
We find that the simplified DAC architecture generally performs on par with the regular one. However, the regular architecture demonstrates superior performance in more challenging domains, particularly dog and assembly tasks. This suggests that the regular DAC approach of defining the optimistic policy as perturbation offers advantages in complex environments. The two architectures differ in their approaches to controlling the KL divergence between the optimistic and pessimistic actors. In the regular DAC setup, the optimistic policy, being a perturbation of the pessimistic policy, can reduce its output to achieve a desired divergence level, independently of the pessimistic policy location. Conversely, in the simplified architecture, the independent optimistic actor must continually adapt to align with the pessimistic policy across the state-action space - posing a challenge in maintaining a stable divergence. To this end, we find that the simplified DAC architecture is not able to meet the target divergence on the dog and assembly task.
@@ D.2 Layer Normalization
Recently, there has been a surge of works exploring the importance of critic regularization in high replay regimes (Liu et al., 2020; Laskin et al., 2020; Hiraoka et al., 2021; Gogianu et al., 2021; Li et al., 2022; D’Oro et al., 2022). Whereas there is still a lot to understand about the interplay of TD learning and network regularization, it is clear that a regularized critic allows a higher replay ratio to be used (Hiraoka et al., 2021; Li et al., 2022; D’Oro et al., 2022). In this paper, we explore only one method that could be considered regularization, that is full-parameter resets (Nikishin et al., 2022; D’Oro et al., 2022). Given the effectiveness of the only regularization we tested, we hypothesize that experimenting with methods like layer normalization (Ba et al., 2016), weight decay (Krogh & Hertz, 1991), spectral normalization (Gogianu et al., 2021) or dropout (Srivastava et al., 2014) could further improve DAC performance. We investigate the effects of layer normalization on DAC performance in Figure 12 and leave further regularization studies for future work.
Figure 12: We investigate the effects of layer normalization on DAC critic and find that it significantly improves performance on the considered locomotion tasks. 10 seeds per task.
We find that using layer normalization on critic within the DAC design significantly improves its performance on locomotion tasks. This results is in line with previous work that augmented other RL algorithms with layer normalization (Li et al., 2022; Hiraoka et al., 2021; Lyle et al., 2023; Ball et al., 2023; Cetin & Celiktutan, 2023)
@@ D.3 Critic Overestimation
$O_{\theta}=\frac{Q_{\theta}(s,a)}{Q^{\pi}(s,a)}$ (49)
@@
Where $O_{\theta}$ denotes the overestimation metric, $Q_{\theta}(s,a)$ is the critic output and $Q^{\pi}(s,a)$ stands for the observed empirical returns. Whereas the metric does not perfectly measure the overestimation, it allows to investigate the relative overestimation in the group of considered algorithms. As shown in Figure 13, we find that the dual architecture prevents the overestimation associated with optimistic TD updates and yields measurements similar to other risk-aware algorithms.
Figure 13: We investigate if the dual architecture allows to mitigate the value overestimation associated with optimistic algorithms. Indeed, the measured overestimation (shown on the $Y$-axis) is similar across all tested algorithms. 10 seeds per task.

### v3

@@ C.1 Pessimistic Actor
The pessimistic actor, denoted as $\pi^{p}_{\phi}$, employs a soft policy target for optimization, as outlined in Equation 8. This actor operates entirely off-policy, utilizing data exclusively from the optimistic actor. The soft policy approach not only promotes state-dependent exploration (Haarnoja et al., 2018a) but also helps in maintaining the hyperbolic tangent (Tanh) output in a non-saturated state (Wang et al., 2020a). Moreover, the pessimistic actor’s inherent non-zero variance contributes to stabilizing the TD learning for the critic. While separating exploration and exploitation theoretically allows for zero variance in TD target sampling, our empirical evidence suggests that introducing noise regularizes the critic (see Table 1).
(a) Normal
(b) Tanh-Normal
(c) Pre-Tanh $\mu$
(d) Post-Tanh Actions
Figure 7: Soft policy learning prevents policy saturation. Figures 7(a) and 7(b) show three Gaussians and the Tanh transformed counterparts. Before Tanh transformation, all policies have equal variance (7(a)). However, since Tanh is nonlinear, applying the transformation changes the policy variance depending on the location of the mean (7(b)). Because of this effect, when using soft policy learning, the agent incurs loss when moving along the saturated portions of the Tanh activation function. Figures 7(c) and 7(d) show policy means and executed actions for SAC with and without soft policy learning on a Humanoid-Stand task. As follows, SAC without the entropy objectives follows a bang-bang policy, whereas regular SAC anchors the policy within the non-saturated portion of Tanh. To this end, soft policy learning reduces the risk of a bang-bang policy.
In the context of representing a policy with a Tanh-Normal distribution, soft policy learning serves two critical functions: promoting entropy and preventing saturation. Firstly, soft policy learning guarantees a certain degree of policy entropy, essential for exploration during the learning process (Haarnoja et al., 2018a). This approach contrasts with methods like TD3 or TOP, where constant entropy is maintained across various states. By optimizing an objective function that balances entropy maximization with Q-values, our policy exhibits variable entropy depending on the state, encouraging exploration particularly in states with lower Q-values. Furthermore, the Tanh function’s inherent saturation characteristic impacts the exploration. The further the unsquashed distribution’s mean is from zero, the more the variance reduces post-squashing, as depicted in Figures 7(a) and 7(b). Implementing soft policy learning thus incurs a loss when the policy deviates from zero. Figures 7(c) and 7(d) illustrate how this approach results in policies predominantly residing within the non-saturated Tanh region (Wang et al., 2020a), effectively avoiding extreme bang-bang policies (Seyde et al., 2021).
@@ C.2 Optimistic Actor
(a) Critic Variance
(b) Pessimistic Actor
(c) Optimistic Actor
Figure 8: Pessimistic underexploration and state-action space coverage on the Pendulum task with state representation embedded into $1$ dimension. The dots represent $500$ state-action samples gathered using the latest policy (pessimistic (black) or optimistic (red)). Figure 8(a) displays the standard deviation ($\sigma$) of the two critics, with smaller values observed in well-explored state-action regions. In Figure 8(b), we depict pessimistic policy probabilities. Due to lower bound optimization, the actor prioritizes state-action subspaces that have already been explored and do not yield critic disagreement. Figure 8(c) illustrates optimistic policy probabilities. Despite having similar entropy levels, following the upper bound policy results in better coverage within critic disagreement regions.
@@
    Reducing the degree of off-policy learning - Given that exploration is conducted solely by the optimistic actor, the pessimistic actor updates entirely off-policy, using transitions sampled from the optimistic policy. This can potentially lead to unstable learning, a phenomenon known as the "deadly triad" (Sutton & Barto, 2018). Integrating KL divergence into the optimistic objective aligns the transitions sampled from the optimistic policy with the expectations under the pessimistic policy, thereby mitigating the extent of off-policy learning.
@@
C.3 PseudoCode
1:  Input: $\pi^{p}_{\phi}$ pessimistic actor; $\pi^{o}_{\eta}$ optimistic actor; $Q_{\theta}$ double critic; $Q_{\bar{\theta}}$ target critic; $\alpha$ temperature; $\beta^{o}$ optimism; $\tau$ KL weight;
2:  Hyperparameters: $\beta^{p}$ pessimism; $\mathcal{KL}^{*}$ target KL;
[…]
5:  for $i=1$ to ReplayRatio do
[…]
7:     Update critic using pessimistic actor (Eq. 9) $\theta\leftarrow\theta-\nabla_{\theta}\bigl{(}Q_{\theta}(s,a)-(r+\gamma($$Q^{\mu}_{\bar{\theta}}(s^{\prime},a^{\prime})+\beta^{p}Q^{\sigma}_{\bar{\theta}}(s^{\prime},a^{\prime})$$-\alpha\log\pi^{p}_{\phi}(a^{\prime}|s^{\prime})\bigr{)}^{2},~{}a^{\prime}\sim\pi^{p}_{\phi}(s^{\prime})$
8:     Update pessimistic actor (Eq. 8) $\phi\leftarrow\phi+\nabla_{\phi}\bigl{(}$$Q^{\mu}_{\theta}(s,a)+\beta^{p}Q^{\sigma}_{\theta}(s,a)$$-\alpha\log\pi^{p}_{\phi}(a|s)\bigr{)}~{}~{}~{}~{}\text{with}~{}~{}~{}~{}a\sim\pi^{p}_{\phi}(s)$
9:     Update optimistic actor (Eq. 7) $\eta\leftarrow\eta+\nabla_{\eta}\bigl{(}Q^{\mu}_{\theta}(s,a)+\beta^{o}Q^{\sigma}_{\theta}(s,a)-\tau KL\bigl{(}\pi_{\phi}^{p}(s)|\bar{\pi}_{\eta}^{o}(s)\bigr{)}\bigr{)}\quad\text{with}\quad a\sim\pi^{o}_{\eta}(s)$
10:     Update entropy temperature $\alpha\leftarrow\alpha-\nabla_{\alpha}\alpha\bigl{(}\mathcal{H}^{*}-\mathcal{H}(s)\bigr{)}$
11:     Update optimism (Eq. 10) $\beta^{o}\leftarrow\beta^{o}-\nabla_{\beta^{o}}(\beta^{o}-\beta^{p})(\frac{1}{|A|}KL(\pi_{\phi}^{p}|\pi_{\eta}^{o})-\mathcal{KL}^{*})$
12:     Update KL weight (Eq. 10) $\tau\leftarrow\tau+\nabla_{\tau}\tau(\frac{1}{|A|}KL(\pi_{\phi}^{p}|\pi_{\eta}^{o})-\mathcal{KL}^{*})$
13:     Update target network $\bar{\theta}\leftarrow\textsc{Polyak}(\theta,\bar{\theta})$
14:  end for
Figure 9: Pseudo-code shows DAC training step, where changes with respect to SAC are colored.
C.4 Optimistic Actor Implementation
Both optimistic and pessimistic policies are modelled using the Tanh-transformed Gaussian distribution denoted as $\mathcal{T}(\mu,\sigma)$, where $\mu$ and $\sigma$ denote the parameters of the underlying Gaussian (Haarnoja et al., 2018a; Wang et al., 2020a). We implement the optimistic actor as a perturbation to the pessimistic policy parameters:
$\begin{split}&\pi_{\eta}^{o}(s)=\mathcal{T}((\mu^{p}_{\phi}(s)+\mu^{o}_{\eta}(s)),(\sigma^{p}_{\phi}(s)*\sigma^{o}_{\eta}(s))).\end{split}$ (24)
Above, $\mu^{p}_{\phi}(s)$ and $\sigma^{p}_{\phi}(s)$ denote mean and standard deviation given by the pessimistic actor network, $\mu^{o}_{\eta}(s)$ and $\sigma^{o}_{\eta}(s)$ denote the optimistic perturbation to the pessimistic policy. Such design allows the optimistic actor network to trivially reduce the KL divergence - to minimize divergence, the optimistic actor network needs to adjust its output towards 00, independently of the values of the pessimistic policy parameters. This design allows the model to keep the KL low, which as we show in Table 1 is crucial for good performance.
C.5 Adjustment of $\beta^{o}$ and $\tau$
@@
Since values of $Q^{\mu}_{\theta}$ and $Q^{\sigma}_{\theta}$ depend on reward scales, as well as aleatoric and epistemic uncertainty of the environment, the value of $\beta^{o}$ cannot be easily set. DAC leverages an observation that for $\beta^{o}=-\beta^{p}$ the optimistic actor recovers the objective of the pessimistic actor. Then, $\beta^{o}$ can be defined such that the divergence between the pessimistic baseline policy and the optimistic policy reaches a desired level, as shown in Equation 10. This adaptive approach, as illustrated in Figure 2, accommodates different scales of Q-values and contrasts with setups like OAC, where optimism is predefined by fixing $\beta^{o}$ at a specific value. We find that the optimism and KL weight used by DAC are highly variable between tasks and even phase of the training, showing the advantage of the automatic adjustment mechanism as opposed to using predefined levels. Furthermore, we find that DAC uses higher levels of optimism than values that were found optimal for OAC (Ciosek et al., 2019).
@@ Appendix D Additional Experiments
D.1 Wallclock performance
Firstly, we assess the wallclock efficiency of DAC as compared to SAC. Figure below investigates the performance of DAC and SAC in both low and high replay regimes.
Figure 10: DAC wall-clock runtime for $1$mln steps is around $15\%$ longer than SAC. Despite additional computational requirements associated with the dual actor, the improvements stemming from DAC lead to a beneficial compute/performance tradeoff. The $Y$-axis denotes IQM with $95$% CI. The $X$-axis denotes the environment steps (left) and wallclock (right) for $10$ seeds. $30$ tasks, $1$mln steps and $10$ seeds. Experiments were conducted on an NVIDIA A100 GPU with 40GB of RAM.
@@ D.2 Layer Normalization
Recently, there has been a surge of works exploring the importance of critic regularization in high replay regimes (Liu et al., 2020; Laskin et al., 2020; Hiraoka et al., 2021; Gogianu et al., 2021; Li et al., 2022; D’Oro et al., 2022). Whereas there is still a lot to understand about the interplay of TD learning and network regularization, it is clear that a regularized critic allows a higher replay ratio to be used (Hiraoka et al., 2021; Li et al., 2022; D’Oro et al., 2022). In this paper, we explore only one method that could be considered regularization, that is full-parameter resets (Nikishin et al., 2022; D’Oro et al., 2022). Given the effectiveness of the only regularization we tested, we hypothesize that experimenting with methods like layer normalization (Ba et al., 2016), weight decay (Krogh & Hertz, 1991), spectral normalization (Gogianu et al., 2021) or dropout (Srivastava et al., 2014) could further improve DAC performance.
Figure 11: We compare high-replay DAC performance with and without layer normalization, as well as the ensemble based, high-replay RedQ (Hiraoka et al., 2021). We use $10$ DMC tasks listed in Table 2 and run $10$ seeds for $1$mln environment steps.
As shown in Figure 11, we find that using layer normalization on critic within the DAC design significantly improves its performance on locomotion tasks. This results is in line with previous work that augmented other RL algorithms with layer normalization (Li et al., 2022; Hiraoka et al., 2021; Lyle et al., 2023; Ball et al., 2023; Cetin & Celiktutan, 2023)
@@ D.3 Critic Overestimation
$O_{\theta}=\frac{|Q_{\theta}(s,a)|}{|G^{\pi}(s,a)|}$ (25)
@@
Where $O_{\theta}$ denotes the overestimation metric, $Q_{\theta}(s,a)$ is the critic output and $G^{\pi}(s,a)$ stands for the observed empirical returns. Whereas the metric does not perfectly measure the overestimation, it allows to investigate the relative overestimation in the group of considered algorithms. As shown in Figure 12, we find that the dual architecture prevents the overestimation associated with optimistic TD updates and yields measurements similar to other risk-aware algorithms.
Figure 12: We investigate overestimation in DAC, SAC and TOP. We use 15 tasks listed in Table 2 and run $10$ seeds for $500$k environment steps.
D.4 Distributional DAC
As discussed in Section 4, our algorithmic implementation follows SAC/TD3 in that it uses a non-distributional double critic, where each critic is taught to output the expected value of returns (Haarnoja et al., 2018a; Fujimoto et al., 2018; Chen et al., 2020; Kostrikov et al., 2020). As such, the critic ensemble disagreement measures the entire uncertainty associated with Q-value approximation. Alternatively, the critic can be implemented to learn the full distribution of returns in an approach referred to as distributional RL Bellemare et al. (2017); Dabney et al. (2018b, a); Moskovitz et al. (2021). In this section, we investigate DAC implementation where the exploration policy is optimistic with respect to aleatoric, epistemic or total uncertainty. Following TOP (Moskovitz et al., 2021) we implement DAC with double quantile distributional critics, where each critic outputs $K$ quantiles of the return distribution (Dabney et al., 2018a). In such setup, the $kth$ quantile of the distribution at state-action $(s,a)$, denoted as $Q^{\pi}_{k}(s,a)$, is learned via optimization of Huber loss of the distributional Bellman error Dabney et al. (2018a) defined as $\delta^{\pi}_{k}=Z^{\pi}(s,a)-Q^{\pi}_{k}(s,a)$ with $Z^{\pi}(s,a)$ representing a sample from the return distribution. We note, that following the Maximum Entropy formulation (Haarnoja et al., 2018a), we define the returns $Z^{\pi}(s,a)$ as a sum of discounted rewards and policy entropy. Furthermore, the point estimate of Q-value is calculated by $Q^{\pi}(s,a)=\frac{1}{K}\sum_{k=1}^{K}Q^{\pi}_{k}(s,a)$. Then, an ensemble of quantile critics can be used to explicitly model both aleatoric and epistemic uncertainties (Moskovitz et al., 2021; Lockwood & Si, 2022). To model aleatoric and epistemic uncertainties, we use the procedure outlined in TOP (Moskovitz et al., 2021). As such, we model aleatoric uncertainty by learning the quantiles of the state-action return distribution, which we denote as $\mathcal{Z}^{\pi}(s,a)$. Thus, we define Q-value as $Q^{\pi}=\mathrm{E}~{}Z^{\pi}(s,a)$ with $Z^{\pi}(s,a)\sim\mathcal{Z}^{\pi}(s,a)$. The spread of the learned distribution $\mathcal{Z}^{\pi}(s,a)$ models the aleatoric uncertainty associated with the policy and the environment. Similarly, we model epistemic uncertainty by the ensemble disagreement when approximating a particular quantile of the return distribution. We consider three implementations of the optimistic policy in a distributional quantile DAC. All approaches apply the optimistic policy optimization as defined in Equation 7, but differ in how the point estimate of Q-value upper-bound is calculated:
Optimistic policy wrt. epistemic uncertainty - In this setup, for a fixed policy $\pi$, we define the Q-value upper bound as $Q^{\beta^{o}}(s,a)=\frac{1}{K}\sum_{k=1}^{K}Q^{o}_{k}(s,a)$ and $Q^{o}_{k}(s,a)=Q^{\mu}_{k}(s,a)+\beta^{o}Q^{\sigma}_{k}(s,a)$, where $Q^{\mu}_{k}$ and $Q^{\sigma}$ represent the ensemble mean and standard deviation in $kth$ quantile. As such, the optimistic policy promotes actions with high epistemic uncertainty measured as critic ensemble disagreement in approximation of the $kth$ quantile.
Optimistic policy wrt. aleatoric uncertainty - Here, for a given policy, we calculate the Q-value upper bound as $Q^{\beta^{o}}(s,a)=\frac{1}{K}\sum_{k=1}^{K}Q^{\mu}_{k}(s,a)+\beta^{o}\sigma_{Z}(s,a)$ where $\sigma_{Z}(s,a)$ is the standard deviation of $K$ quantile outputs, averaged over the critic ensemble. As such, the optimistic policy promotes actions with high aleatoric uncertainty measured as the spread of the return distribution.
Optimistic policy wrt. both uncertainties - Finally, we consider an optimistic policy that promotes both aleatoric and epistemic uncertainties, albeit still explicitly modeled via an ensemble of quantile critics. Here, for a given policy we calculate the upper-bound Q-value by $Q^{\beta^{o}}(s,a)=\frac{1}{K}\sum_{k=1}^{K}Q^{o}_{k}(s,a)+\beta^{o}\sigma_{Z}(s,a)$, with $Q^{o}_{k}(s,a)=Q^{\mu}_{k}(s,a)+\beta^{o}Q^{\sigma}_{k}(s,a)$. As such, the optimistic policy promotes actions that are associated with both epistemic and aleatoric uncertainty.
Figure 13: We run three variations of distributional DAC ($RR=2$) on $10$ tasks from DMC and MW benchmarks. $Y$-axis reports IQM and $X$-axis reports environment steps. 10 seeds per task.
Figure 14: We run three variations of distributional DAC ($RR=16$) on $10$ tasks from DMC and MW benchmarks. $Y$-axis reports IQM and $X$-axis reports environment steps. 10 seeds per task.
We run the three variations of Distributional Quantile DAC for $1$mln environment steps and both replay regimes. The results are presented in the Figures below. We find that all variations achieve similar performance, with the standard implementation promoting the epistemic uncertainty slightly outperforming other variations.